OpenStudy (anonymous):

What is the solution set for the open sentence with the given replacement set? y2 – 3 = y + 9, {0, 1, 3, 4}

OpenStudy (anonymous):

this question then 2 more

OpenStudy (joshawad):

so is y squared or is it 2Y?

OpenStudy (anonymous):

not sure its y2 it might be

OpenStudy (joshawad):

i cant help u if ur not sure

OpenStudy (anonymous):

ill find next question

OpenStudy (anonymous):

Determine the width of the garden. Nina must put 60 feet of fencing around her rectangular garden. The length of the garden is three feet longer than twice the width. What is the solution set for 2w + 2(2w + 3) = 60 given the replacement set {6, 7, 8, 9}?

OpenStudy (joshawad):

so first we distribute 2

OpenStudy (joshawad):

2W + 4W + 6 = 60 combine like kind 6W + 6 -6 = 60 -6 Getting rid of 6 6W = 54 divide by 6 W=9

OpenStudy (anonymous):

ok and 1 more

OpenStudy (anonymous):

Determine the rate at which Philip traveled during the second part of his trip. Philip drives for 4 hours at a rate of 50 miles per hour. Later, he drives for another 3 hours so that his total distance traveled for the day is 365 miles. What is the solution set for 4(50) + 3r = 365 given the replacement set {52, 53, 54, 55}?

OpenStudy (joshawad):

4(50) + 3R = 365 combine like kinds 200 + 3R = 365 taking away 200 200 - 200 + 3R = 365 -200 3R = 165 divide by 3 R =55